Jefferson Lab HALL A Logbook for February 12 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 362572 02/29/12 23:57 mhuang New reconstruction matrix in DB vdc files
  • 362571 02/29/12 23:23 R. Zielinski RHRS RT Module working as expected
  • 362570 02/29/12 23:16 R. Zielinski ROC 2 Working (for now)
  • 362543 02/29/12 16:08 R. Zielinski ROC 2 Error (again)
  • 362505 02/29/12 11:47 M. Cummings Fastbus Crate Resets
  • Tuesday

  • 362476 02/28/12 20:14 Kalyan taking cosmics with ThirdArm
  • 362471 02/28/12 20:02 mhuang Quads tuning numbers
  • 362439 02/28/12 17:34 pengjia fixed happex crash in low rate
  • 362418 02/28/12 11:44 pzhu happex crate crash in low rate
  • 362409 02/28/12 10:22 Kalyan Rough draft of commissioning plan
  • Monday

  • 362408 02/27/12 22:33 Jixie Zhang, G2p Septa field measurement
  • 362407 02/27/12 22:30 T. Badman [Target] Target Magnet Ramping down
  • 362406 02/27/12 21:49 Jixie Zhang G2p Target field measurement
  • 362405 02/27/12 18:57 T. Badman [Target] Magnet Ramping Up
  • 362404 02/27/12 18:07 adaq HALOG test
  • Saturday

  • 362370 02/25/12 13:45 V. Sulkosky VDC HVs Ramped Up
  • 362366 02/25/12 11:04 pengjia Ed Folts don't allow us to go into the Hall
  • 362365 02/25/12 10:55 pzhu BPM receiver installed
  • Friday

  • 362364 02/24/12 18:42 V. Sulkosky Ramped VDC HV's Down
  • 362357 02/24/12 15:09 kelleher TECHS Detectors Gas System Interlock Alarm
  • 362353 02/24/12 13:33 R. Zielinski Happex also rebooting during run on LHRS
  • 362352 02/24/12 13:31 R. Zielinski RHRS Happex Crate Rebooting during run
  • 362346 02/24/12 12:54 R. Zielinski Psuedo High-Rate DAQ test
  • Thursday

  • 362319 02/23/12 16:10 V. Sulkosky/ Fastbus Crate Resets
  • Wednesday

  • 362306 02/22/12 19:50 jmaxwell [Target] Magnet at Full Field
  • 362305 02/22/12 18:28 jmaxwell [Target] Material in Insert
  • 362304 02/22/12 18:16 jmaxwell [Target] Magnet Ramping Up
  • 362287 02/22/12 10:08 kelleher RC report -- Finished version
  • 362286 02/22/12 10:00 kelleher RC report
  • 362285 02/22/12 00:35 jmaxwell [Target] Target Magnet Ramped Down
  • Tuesday

  • 362283 02/21/12 22:22 A. Camsonne DHCP and tftp server setup on adaql10
  • 362282 02/21/12 19:03 Chao Third Arm TS crate number move to ROC 12
  • 362273 02/21/12 17:49 jmaxwell [Target] Ramping Magnet Up
  • 362267 02/21/12 14:43 jmaxwell [Target] Test Post Again
  • 362266 02/21/12 11:04 jmaxwell [Target] Test Entry, Ignore
  • 362265 02/21/12 10:34 Fansler N2 valves
  • 362264 02/21/12 10:33 Fansler/Butl Lead flows
  • 362263 02/21/12 01:01 jmaxwell [Target] Target Magnet Ramped Down
  • Monday

  • 362259 02/20/12 20:30 jmaxwell [Target] Magnetic Field at 5T, Ammonia in Nose
  • 362256 02/20/12 17:51 brads Beamline HV GUI working again
  • Saturday

  • 362229 02/18/12 19:02 brads re: Beamline HV crate control not working (late entry)
  • 362228 02/18/12 12:49 R. Zielinski Start of run 21514 -- RHRS_ROC6 Configuration
  • 362226 02/18/12 10:39 R. Zielinski Stopping run 21257 - ROC 1 not getting gates
  • 362225 02/18/12 08:49 A. Camsonne FB error file
  • 362224 02/18/12 08:48 A. Camsonne FB error file
  • 362223 02/18/12 07:47 A. Camsonne Started run 21527 RHRS_ROC6 configuration
  • 362222 02/18/12 06:15 A. Camsonne Started run 21526 RHRS_ROC6 configuration
  • 362221 02/18/12 01:34 R. Zielinski RHRS ROC 2 Problems
  • 362220 02/18/12 00:42 M. Cummings HRS DAQ Update
  • Friday

  • 362212 02/17/12 16:21 A. Camsonne Took back spare helicity board
  • Thursday

  • 362205 02/16/12 21:07 brads Beamline HV crate control not working
  • Wednesday

  • 362197 02/15/12 22:04 Jack Left Arm Fastbus cabling
  • 362196 02/15/12 15:54 A. Camsonne Compton electron detector vacuum pumped dow.
  • Tuesday

  • 362191 02/14/12 21:00 R. Zielinski RHRS DAQ Update
  • Monday

  • 362171 02/13/12 18:54 V. Sulkosky SFI and Fastbus status on Right Arm
  • 362169 02/13/12 17:30 A. Camsonne Replaced fastbus boards on ROC 5
  • 362168 02/13/12 17:04 A. Camsonne SFI and Fastbus status on Left Arm
  • Thursday

  • 362167 02/09/12 18:58 V. Sulkosky hareboot32 Added to RHRS for Happex Crate
  • Wednesday

  • 362162 02/08/12 10:44 Chao Scaler ring buffer readout has been separated from scaler count readout
  • 362161 02/08/12 10:17 slifer status read back for He bags (from Heidi)
  • 362160 02/08/12 09:10 Pengjia bcm signal split problem
  • Tuesday

  • 362157 02/07/12 20:01 Jack Left Arm Fastbus turned off
  • 362154 02/07/12 17:57 V. Sulkosky hareboot28 Moved to LHRS for Happex Crate
  • Monday

  • 362144 02/06/12 17:53 A. Camsonne cptsv1 portserver dead
  • 362130 02/06/12 17:13 romanip VME CPU for Moller FADC DAQ has been replaced
  • Saturday

    
    

    Friday

    
    

    Wednesday

  • 362046 02/01/12 11:05 Sasha Artificial beam charge asymmetry test